The radiator in a car is filled with a solution of 70% antifreeze and 30% water. The manufacturer of the antifreeze suggests that, for summer driving, optimal cooling of the engine is obtained with only 50% antifreeze. If the capacity of the radiator is 3.5 L, how much coolant should be drained and replaced with water to reduce the antifreeze concentration to the recommended level?
Answer by stanbon(75887)

The radiator in a car is filled with a solution of 70% antifreeze and 30% water. The manufacturer of the antifreeze suggests that, for summer driving, optimal cooling of the engine is obtained with only 50% antifreeze. If the capacity of the radiator is 3.5 L, how much coolant should be drained and replaced with water to reduce the antifreeze concentration to the recommended level?
----
Equation:
water - water + water = water
0.30*3.5 - 0.30x + 1*x = 0.5*3.5
----
Multiply thru by 100 to get:
30*3.5 - 30x + 100x = 50*3.5
70x = 20*3.5
---
x = 1 liter (amt. of solution that should be replaced)
==============
Cheers,
Stan H.